# YOLO Evolution: A Comprehensive Benchmark and Architectural Review of YOLOv12, YOLO11, and Their Previous Versions

### v3

@@ V Conclusion
This benchmark study thoroughly evaluates the performance of various YOLO algorithms. It pioneers a comprehensive comparison of YOLO11 against its predecessors, evaluating their performance across three diverse datasets: Traffic Signs, African Wildlife, and Ships and Vessels. The datasets were carefully selected to encompass a wide range of object properties, including varying object sizes, aspect ratios, and object densities. We showcase the strengths and weaknesses of each YOLO version and family by examining a wide range of metrics such as Precision, Recall, Mean Average Precision (mAP), Processing Time, GFLOPs count, and Model Size. Our study addresses the following key research questions:
@@
  • What are the specific strengths and limitations of each YOLO version, and how can these insights inform the selection of suitable algorithms for various applications?
@@
YOLOv12, despite being the newest addition to the YOLO family, delivered underwhelming performance. Its ambitious architectural changes, including the Area Attention Module and R-ELAN, introduced complexity without translating into clear advantages. The model struggled to balance accuracy and speed, highlighting the challenges of integrating advanced attention mechanisms effectively within the YOLO framework.

### v4

@@ 7 Conclusion
This benchmark study thoroughly evaluates the performance of various YOLO algorithms. It pioneers a comprehensive comparison of YOLOv12 against its predecessors and assesses their performance across three diverse datasets: Traffic Signs, African Wildlife, and Ships and Vessels. The datasets were carefully selected to encompass various object properties, including varying object sizes, aspect ratios, and object densities. We showcase the strengths and weaknesses of each YOLO version and family by examining a wide range of metrics such as Precision, Recall, Mean Average Precision (mAP), Processing Time, GFLOPs count, and Model Size. Our study addresses the following key research questions:
  • 1.
@@
  • 3.
    What are each YOLO version’s specific strengths and limitations, and how can these insights inform the selection of suitable algorithms for various applications?
  • 4.
    Which YOLO models are best suited for specific real-world scenarios?
@@
Despite being the newest addition to the YOLO family, YOLOv12 delivered an underwhelming performance. Its ambitious architectural changes, including the Area Attention Module and R-ELAN, introduced complexity without translating into clear advantages. The model struggled to balance accuracy and speed, highlighting the challenges of effectively integrating advanced attention mechanisms within the YOLO framework.
@@
Declarations
  • 1.
    Funding: Not applicable
  • 2.
    Conflict of interest/Competing interests: The authors declare no conflict of interest.
  • 3.
    Ethics approval and consent to participate: Not applicable
  • 4.
    Consent for publication: The authors confirm that all necessary permissions have been obtained for the publication of this work, including the use of copyrighted material and any personal information or sensitive data.
  • 5.
    Data availability: Data is available on GitHub and will be publicly available upon publication of the final version
  • 6.
    Materials availability: Not applicable
  • 7.
    Code availability: Source code is available on GitHub and will be publicly available upon publication of the final version.
  • 8.
    Author contribution: All authors contributed to the study conception and design. Material preparation, data collection, and analysis were performed by Nidhal Jegham and Chan Young Koh. The first draft of the manuscript was written by Nidhal Jegham, and all authors commented on and edited previous versions of the manuscript. All authors read and approved the final manuscript. Dr. Abdeltawab Hendawi supervised the work of this study.